What is the result of graphing the equation (x − h)2 + (y − k)2 = r2 ?
Answer by Alan3354(69443)

What is the result of graphing the equation (x − h)2 + (y − k)2 = r2 ?
-----
Use ^, Shift 6, for exponents.
==================
(x − h)^2 + (y − k)^2 = r^2 is a circle.
Center at (h,k) and radius r.
======================
DL the FREE graph software at
www.padowan.dk
----
Use F6, enter, eg, (x − 1)^2 + (y − 2)^2 = 3^2